If 2000 cm3 of 4% alcohol was combined with 400 cm3 of 10% alcohol, how much of a solution of 50% alcohol should be added to make a solution of 20% alcohol?
Answer by Alan3354(69443)

If 2000 cm3 of 4% alcohol was combined with 400 cm3 of 10% alcohol, how much of a solution of 50% alcohol should be added to make a solution of 20% alcohol?
----------------
Add C cc's of 50%
---
2000*4 + 400*10 + C*50 = (2000+400+C)*20 = 48000 + 20C
12000 + 50C = 48000 + 20C
30C = 36000
C = 1200 cc's of 50%